the sum of two numbers is 97. the difference of the two numbers is 57. find the two numbers
Answer by jonvaliente(64)

Let x=first number and y=second number
x+y=97
x-y=57
.
If we add these two equations we get:
2x=154
x=77
since x+y=97, then:
77+y=97
y=97-77
y=20
The two numbers are 77 and 20